OpenStudy (highschoolmom2010):

A design on the surface of a balloon is 5 cm wide when the balloon holds 71 cm^3 of air. How much air does the balloon hold when the design is 10 cm wide? Explain the method you use to find the amount of air.

OpenStudy (anonymous):

I believe this is the volume of a sphere formula, are you familiar with it?

OpenStudy (highschoolmom2010):

not really

OpenStudy (agent0smith):

Hmm... this isn't super straightforward. I think first things first, find the radius.I have an idea how to do this, but it seems like a fair bit of work, maybe there's an easier way. \[\Large V = \frac{ 4 }{ 3 } \pi r^3\]you're given volume = 71 so plug that in, find r.

OpenStudy (highschoolmom2010):

\[71 cm^3=\frac{ 4 }{ 3 }\pi r^2\]

OpenStudy (highschoolmom2010):

thats as far as i got....

OpenStudy (highschoolmom2010):

ok let me do that

OpenStudy (agent0smith):

@satellite73 maybe you or someone else can think of an easier way to solve this than what i have in mind :/ maybe we can use the surface area, instead of needing things like arc length.

OpenStudy (agent0smith):

Maybe we can use the proportion of surface area to volume to make this easier...

OpenStudy (highschoolmom2010):

\[71*3=\frac{ 4 }{ 3 }\pi r ^2*3\] \[\frac{ 123 }{ 4 }=\frac{ 4\pi r^2 }{ 4 }\] \[\frac{ 30.75 }{ \pi }=\frac{ \pi r^2 }{ \pi }\] \[9.788=r^2\] am i on the right track

OpenStudy (agent0smith):

I think so, but maybe we don't even need the radius... but i guess it might come in handy. maybe we can use the fact that the surface area increases by a factor (of 4 if i'm thinking about it right) so the volume must also increase by a constant factor - which would make this much simpler. @terenzreignz

OpenStudy (agent0smith):

Like if the surface area quadruples (which i *think* it does based on the design radius doubling...), that'd mean the radius must have doubled - which would mean the volume must be octuple what it was. I'm just not sure this is actually legit.

OpenStudy (agent0smith):

@e.mccormick

OpenStudy (agent0smith):

@shamil98

OpenStudy (highschoolmom2010):

@thomaster @radar

OpenStudy (anonymous):

is this some kind of proportionality problem? (5/10)^3 = 71/V, and solve for V

OpenStudy (highschoolmom2010):

i dont think so not sure to be honest

OpenStudy (anonymous):

V = 568, a whole number, so it must be right XD But then still, that is based on the assumption that the length of the design grows proportionally as the balloon gets bigger. Of course, i'm not sure if that is actually true in real life

OpenStudy (e.mccormick):

\(S\cdot\dfrac{r}{3}=V\)

OpenStudy (highschoolmom2010):

@e.mccormick can you explain a little @sourwing im lost on how you got 568

OpenStudy (e.mccormick):

And you can find the radius from the first volume. Since your looking at just the width you are only changing in one dimension.

OpenStudy (anonymous):

(5/10)^3 = 71/V V(5/10)^3 = 71 V = 71 / (5/10)^3 = 568

OpenStudy (highschoolmom2010):

(5/10)^3 would i need to reduce that?

OpenStudy (anonymous):

if you want to. The final number is the same anyway

OpenStudy (e.mccormick):

Yep, and it checks against doing it with more steps.

OpenStudy (highschoolmom2010):